# DDaTR: Dynamic Difference-aware Temporal Residual Network for Longitudinal Radiology Report Generation

### v1

@@ IV-C Ablation Study and Case Study
Effect of DAM and FAM. DAM and FAM represent the configurations where dynamic fusion is removed from the DDAM and DFAM modules, respectively. In our ablation studies, we separately removed DAM and FAM to assess their impacts on the model’s performance. As indicated in Tab. III, the FAM led to an F1 increase of 0.025 on the MIMIC-CXR and 0.011 on the Longitudinal-MIMIC. These results demonstrate the effectiveness of FAM in extracting prior visual features that are better aligned with clinical semantics. Similarly, DAM resulted in an an increase of 0.039 in the F1 score on the MIMIC-CXR and 0.018 on the Longitudinal-MIMIC, demonstrating its effectiveness in enhancing the model’s ability to capture temporal changes across visits. The comparison between the two sets of experiments indicates that the DAM plays a more critical role, with substantial impact on both the CE and NLG metrics. Moreover, the FAM, which introduces clinical intent through the prior report, shows a notable improvement in CE but has limited effect on NLG. This suggests that the prior report primarily contributes to enhancing diagnostic accuracy.
@@
Effect of dynamic fusion. Subsequently, we evaluated the effect of the dynamic fusion mechanism itself, which adaptively integrates longitudinal features. As shown in Tab. III, the overall improvement in F1 score is primarily attributed to an increase in precision, while the recall remains unchanged. We attribute this to the effectiveness of the dynamic fusion mechanism that enables the model to better handle scenarios where prior information may be missing or noisy, thereby improving diagnostic precision.
@@
Case study. Fig.8 presents a comparison of generated results. As can be observed, our method accurately captures the changes between two period images and provides precise descriptions. Notably, in case A, it can clearly articulate detailed descriptions, such as ‘one-third of the left hemithorax’. We analyzed the training data and found that around 10 cases contain this type of description, such as “the effusion occupies approximately one-third of the left hemithorax”, “occupying approximately one-third of the left hemithorax”. Although the occurrence is relatively rare, our method appears to have captured the associated patterns and is able to generate similar expressions during testing when encountering similar cases. In Case B, although our generated report is more accurate than those of the baselines, the ground truth indicates uncertainty regarding effusions, while all models including ours make a definitive statement. This highlights an ongoing challenge in distinguishing between definite and probable diagnoses, which can vary due to differing interpretations of uncertainty among radiologists with different levels of experience. In Case C, the model occasionally generates additional negative findings or extra information, such as ”mediastinal clips”.
@@
Visualization. To better demonstrate how the model captures and utilizes longitudinal information, we conducted the visualization experiments of our DDATR. Specifically, we generated Grad-CAM [64] visualizations for both the current and previous images according to diagnostic labels, as shown in Fig. 9. The target layer is the final normalization layer of our DDATR. We present ten examples, analyzing different diagnostic categories and progressions.
@@ V Conclusion
However, our model still has certain limitations that require further investigation and improvement. First, the diagnostic accuracy of current report generation models including ours still does not reach the level of a real radiologist. Therefore, in clinical practice, such models are primarily used to assist radiologists in improving efficiency, with the generated reports still requiring manual review and correction. Secondly, although our method has comprehensive experiments on the X-ray report generation task, its effectiveness on other imaging modalities, such as ultrasound or CT, remains to be further explored and analyzed. Third, each component of our longitudinal framework still holds potential for further optimization, which could lead to further performance gains in future work.

### v2

@@ 4.3 Ablation Study and Case Study
Effect of DAM and FAM. DAM and FAM represent the configurations where dynamic fusion is removed from the DDAM and DFAM modules, respectively. In our ablation studies, we separately removed DAM and FAM to assess their impacts on the model’s performance. As indicated in Tab. 3, the FAM led to an F1 increase of 0.025 on the MIMIC-CXR and 0.011 on the Longitudinal-MIMIC. These results demonstrate the effectiveness of FAM in extracting prior visual features that are better aligned with clinical semantics. Similarly, DAM resulted in an an increase of 0.039 in the F1 score on the MIMIC-CXR and 0.018 on the Longitudinal-MIMIC, demonstrating its effectiveness in enhancing the model’s ability to capture temporal changes. The comparison between the two experiments indicates that the DAM plays a more critical role, with substantial impact on both the CE and NLG metrics. Moreover, the FAM, which introduces clinical intent of the prior report, shows a notable improvement in CE but has limited effect on NLG. This suggests that the prior report primarily contributes to enhancing diagnostic accuracy.
@@
Effect of dynamic fusion. We evaluated the effect of the dynamic fusion mechanism itself, which adaptively integrates features. As shown in Tab. 3, the overall improvement in F1 is primarily attributed to an increase in precision, while the recall remains unchanged. We attribute this to the effectiveness of the dynamic fusion mechanism that enables the model to better handle scenarios where prior information may be missing or noisy, thereby improving diagnostic precision.
@@
Case study. Fig.8 presents a comparison of generated results. As can be observed, our method accurately captures the changes between two period images and provides precise descriptions. Notably, in case A, it can clearly articulate detailed descriptions, such as ‘one-third of the left hemithorax’. We analyzed the training data and found that around 10 cases contain this type of description, such as “the effusion occupies approximately one-third of the left hemithorax”, “occupying approximately one-third of the left hemithorax”. Although the occurrence is relatively rare, our method appears to have captured the associated patterns and is able to generate similar expressions during testing when encountering similar cases. In Case B, we observed that the generated reports still exhibit misalignment issues. Although ours is more accurate than those of the baselines, the ground truth indicates uncertainty regarding effusions, while all models including ours make a definitive statement. This highlights an ongoing challenge in distinguishing between definite and probable diagnoses, which can vary due to differing interpretations of uncertainty among radiologists with different levels of experience. In Case C, we observed that the generated reports still exhibit hallucinations issues. The model occasionally generates additional negative findings or extra information, such as ”mediastinal clips”. In addition, there are still many cases with diagnostic errors reflecting omissions or false positive issues.
@@
Visualization. To better demonstrate how the model captures and utilizes longitudinal information, we conducted the visualization experiments of our DDATR. Specifically, we generated Grad-CAM [65] visualizations for both the current and previous images according to diagnostic labels, as shown in Fig. 9. We present ten examples, analyzing different diagnostic categories and progressions.
@@ 5 Conclusion
However, our model still has certain limitations that require further investigation and improvement. First, the diagnostic accuracy of current report generation models including ours still does not reach the level of a real radiologist. Issues such as omissions, hallucinations, and misalignments still persist in the generated reports. Therefore, in clinical practice, such models are primarily used to assist radiologists in improving efficiency, with the generated reports still requiring manual review and correction. Secondly, although our method has comprehensive experiments on the X-ray report generation task, its effectiveness on other imaging modalities, such as ultrasound or CT, remains to be further explored and analyzed. Third, each component of our longitudinal framework still holds potential for further optimization, which could lead to further performance gains in future work.
[…]